DontGo:

The graph will be in the comments (05.03)Calysta sells used video games. In addition to a fixed salary, she earns a commission for each video game she sells. The table shows Calysta’s total earnings, y, (in dollars), from selling x video games: Which equation best shows the relationship between x and y? y = x + 8 y = 8x + 70 y = 8x + 78 y = x + 78

DontGo:

DontGo:

@Kamauri

DontGo:

Can you help me?? I am terrible at math and soo.........

Kamauri:

me to sometimes

DontGo:

the chart will be in a file (05.01)Elliot has been running a lawn care business since 2000. He cuts grass, trims, and weed whacks yards for his customers throughout the season. Each year, he has increased his fee by the same amount. The table shows what Elliot charged each customer for two given years of his business: A. What is the rate of change and initial value for Elliot’s business? How do you know? B. Write an equation in slope-intercept form to represent the fees that Elliot charges each year.

DontGo:

I know the initial value

beholder:

(05.03) Since this is linear, we have \(y = mx+b\) for some constants \(m,b\). From the table, we can find the slope to be \[m = \frac{86-78}{2-1} = 8 \] and we must have \(b = 30\) otherwise when \(x = 1\) we won't have \(y = 38\). The answer is \(y = 8x + 30\).

beholder:

Sorry, \(b = 70\) so answer is \(y = 8x + 70\)

beholder:

(05.01) (A.) The rate of change can be calculated using the slope formula:\[ m = \frac{y_2 - y_1}{x_2-x_1} = \frac{1350 - 750}{2010-2000} = \frac{600}{10} = \$60\text{ per year}.\]. The rate of change is $60 per year. Since the business started in 2000, the initial value is clearly $750 as that is the lawn care fee when he started the business. (B.) The slope-intercept form of the fees is of the form \(y = mx+b\). Here, \(m = 60\) is our rate of change and \(b = 750\) is our initial value / x-intercept. So the slope-intercept form is \(y = 60x + 750\).